# Deep Reasoning Translation via Reinforcement Learning

### v1

@@ 4 Failure Experience
In this section, we discuss our failures in reward design. These experiences could also be regards as the ablation study in $r_{\text{all}}$.
4.1 The Length Penalty of Thought Process
In the pilot RL experiments, we also design a length penalty for the thought content. Specifically, it could be defined as follows:
$r_{\text{length}}=\begin{cases}-\frac{\operatorname{length}(\text{think})-\beta}{\eta}&\text{if}\operatorname{length}\text{(think)}\geq\beta\\ 0&\text{if}\operatorname{length}\text{(think)}<\beta\end{cases}$ (9)
[…]
$r_{\text{all (w/ lp)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}+\alpha\times(r_{\text{thought}}+r_{\text{length}})&\text{if }r_{\text{format}}\neq 0\end{cases}$ (10)
We set $\beta$ to 400, $\eta$ to 400, and $\alpha$ to 20 in the pilot experiments. The DeepTrans-7B model trained with $r_{\text{all (w/ lp)}}$ is denotes as DeepTrans ($r_{\text{all (w/ lp)}}$). As shown in Figure 3, we find that the length penalty does not bring additional improvement to model performance during training. In contrast, it reduces the rate of improvement in model performance. Therefore, we do not adopt the length penalty in the final overall reward.
@@ 4.2 The Effect of Thought Reward
We also attempt to remove the thought reward during RL training, and the variant overall reward is defined as:
$r_{\text{all (w/o th)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}&\text{if }r_{\text{format}}\neq 0\end{cases}$ (11)
However, during RL training, we find that DeepTrans ($r_{\text{all (w/o th)}}$) quickly discards the thought process, and tends to directly output the final translation until the end of the training.
To make the model provide a valuable thought process, we also consider a short length penalty on the thought process. In detail, if the thought length is less than $k$ tokens, we give a huge penalty to the overall reward:
$r_{\text{all (w/o th \& w/ sp)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ 0&\text{elif }\operatorname{length}\text{(think)}<10\\ r_{\text{trans}}&\text{else}\end{cases}$ (12)
In this way, the final trained model tends to only provide several empty words in the thought content, e.g., “Well, first I have to understand the meaning of the sentence and make sure the translation is accurate. Then, I will express it in fluent Chinese, keeping the sentence coherent and natural”. This kind of thinking process loses the relevance and specificity towards source sentences.
Besides, the model trained with $r_{\text{all (w/o th)}}$ or $r_{\text{all (w/o th \& w/ sp)}}$ significantly underperforms DeepTrans-7B in terms of all metrics, and also shows code-mix results in the translations. These findings demonstrate the importance of the thought process in model application, and we should consider how to guide models to generate valuable thought during RL training.
@@
TR. w/ 3-point scale TR. w/ 5-point scale TR. w/ 3-point scale TR. w/ 5-point scale
The man from the West unfolded the little piece of paper handed him. His hand was steady when he began to read, but it trembled a little by the time he had finished. The note was rather short. {CJK}UTF8gbsn 西方人展开递给他的一小张纸条。他开始阅读时手很稳,但读完后微微发抖。那张纸条很短。 [CometKiwi: 80.10] {CJK}UTF8gbsn 那自西而来之人展开了递与他之小小书笺,启阅时手心尚稳,及至终了,已微颤难禁。此简不过寥寥数语[CometKiwi: 66.25] {CJK}UTF8gbsn那自西方而来的男子轻轻展开递至掌心的小小纸笺初时指尖稳健如常,逐字细阅其上墨迹,及至文末,方觉指间微颤不宁。此纸不过寥寥数行,却藏无限意味[CometKiwi: 68.79] {CJK}UTF8gbsn 西边来人展信细阅初启笔端神定如磐,及至文终,指尖微颤。简札不过寥寥数行[CometKiwi: 74.02] {CJK}UTF8gbsn 西边来的人,缓缓展开递到手里的小小纸片。起初,他的双手握得十分稳当,可待读罢之时,指尖已微微有些发颤了。那张纸条实在不算长。 [CometKiwi: 79.41]
When it was light enough Johnsy, the merciless, commanded that the shade be raised. {CJK}UTF8gbsn 当光线足够明亮时,无情的约翰西斯命令把窗帘拉上。 [CometKiwi: 81.72] {CJK}UTF8gbsn 当光线已然充足,约翰西,那无情之人,竟令帷幕徐徐卷起[CometKiwi: 74.46] {CJK}UTF8gbsn 待天色渐明,那冷漠无情的约翰西,便命人将帘幕徐徐卷起[CometKiwi: 75.08] {CJK}UTF8gbsn 天色稍显明亮,约翰西,那颗铁石心肠,便下令将阴影卷起。 [CometKiwi: 66.96] {CJK}UTF8gbsn 待天色足够明亮,约翰西这颗铁石心肠,已然下令将帘幕卷起。 [CometKiwi: 71.40]
When one loves one’s Art no service seems too hard. {CJK}UTF8gbsn 当一个人热爱自己的艺术时,任何服务都显得不值一提。 [CometKiwi: 70.98] {CJK}UTF8gbsn 若钟情于艺道,任劳任怨亦无妨[CometKiwi: 69.38] {CJK}UTF8gbsn 一旦钟爱所钟之艺便觉任劳任怨非艰[CometKiwi: 69.38] {CJK}UTF8gbsn 艺海倾心无难事,任凭千般苦亦甘[CometKiwi: 54.76] {CJK}UTF8gbsn 只要钟爱所钟情的艺术,任何付出都觉轻而易举。 [CometKiwi: 76.59]
For two or three days she went out canvassing for pupils. One evening she came home elated. {CJK}UTF8gbsn 连续两三天她出去招生,一个傍晚她兴高采烈地回家了。 [CometKiwi: 83.55] {CJK}UTF8gbsn 连续几日外出寻觅门徒一夕归家心欢悦[CometKiwi: 80.25] {CJK}UTF8gbsn 她连着几日四处奔走寻觅门徒,直至某晚,她满心欢喜地踏入了家园。 [CometKiwi: 74.54] {CJK}UTF8gbsn 连续几日奔走寻门徒,直至傍晚方归家满心欢喜步履轻,笑颜如花映晚霞[CometKiwi: 55.10] {CJK}UTF8gbsn 她接连走了两三天,到处奔走寻觅学生。一天傍晚,她满面春风地回到了家中。 [CometKiwi: 83.87]
Table 5: Case Studies on O. Henry (TR.: Translation Reward). Purple indicates the translations are in strong Chinese classical literary style.
@@ 4.3 The Effect of Translation Reward
We define the translation reward with a 100-point scale in § 2.1, and we further study the effect of the scoring criteria in translation rewards. Specifically, we use two variants based on the original $r_{\text{trans}}$: $r_{\text{trans (3 point)}}$ and $r_{\text{trans (5 point)}}$ simply narrow the scoring scope of $r_{\text{trans}}$ from a 100-point to a 3-point scale and a 5-point scale, respectively. During training with these two variants, the $\alpha$ in Eq. 4 is also changed to ensure the same ratio between thought rewards and translation rewards. We denote the models trained with the variants as DeepTrans (trans reward w/ 3-point scale) and DeepTrans (trans reward w/ 5-point scale), respectively. As shown in Figure 4 (a), both variants show the sub-optimal results in terms of GRF, verifying the rationality of our original design.
@@
As for CometKiwi scores in Figure 4 (b), we find that DeepTrans (trans reward w/ 5-point scale) shows similar trends with original DeepTrans, i.e., first drops and finally recovers. However, DeepTrans (trans reward w/ 3-point scale) continues to decline and fluctuate during the training progress. In the end, it achieves extremely low performance in terms of CometKiwi, and perhaps it needs more training steps to recover the performance. In view of the stylization phenomenon we discussed in § 3.5 and Table 4, we conjecture the trends shown in Figure 4 (b) are attributed to stylization.
To figure it out, we provide case studies on the variant models in Table 5 (where the cases are the same as Table 4). As we can see, when the models are trained with 300 steps, both variants show the stylization phenomenon. Further, when trained with 600 steps, we find that DeepTrans (trans reward w/ 5-point scale) discards the classical-style translations. In contrast, DeepTrans (trans reward w/ 3-point scale) enhances the classical-style translations, and the results are like ancient poetry.
@@
As we also described in § 3.5, the basic reason of the stylization phenomenon is a bias/preference of the reward model, which encourages the translation model to provide stylized translations as a local optimal solution. When the reward model provides a wider scoring scope (e.g., the original $r_{\text{trans}}$ with 100-point scale), it gives a larger latent space for the policy model to explore. Therefore, the policy model is more likely to get out of the local optimal solution. If using a narrow reward scope, e.g., $r_{\text{trans (3 point)}}$, some minor changes in the policy model will not be reflected by the reward signal immediately, and thus make it hard to transfer the translation strategies. In conclusion, the scoring scope of the reward models has a great effect on the policy model, and it is necessary to employ a wider reward scope for the model to explore a good translation strategy.
5 Related Work
5.1 Deep Reasoning LLMs
In recent years, deep reasoning LLMs, e.g., OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), have pioneered a growing research in long chain-of-thought (CoT) reasoning. Many studies make their efforts on different tasks or scenarios Zhang et al. (2024); Chen et al. (2025b); Li et al. (2025); Guan et al. (2025); Jin et al. (2025). Chen et al. (2025b) provide a systematic survey on deep reasoning LLMs to introduce the development process. Some researchers investigate the MT capability of deep reasoning LLMs. Zhao et al. (2024) and Liu et al. (2025) briefly show that the long CoT reasoning helps the model to reach more idiomatic translations. DRT(-o1) Wang et al. (2024) is further proposed in literature MT, and it is trained with synthesized thought processes as well as translations. An empirical study Chen et al. (2025a) also discusses how well deep reasoning LLMs perform on MT tasks. R1-T1 He et al. (2025) leverages COMET scores as reward signals to train deep reasoning MT models with RL. However, the deep reasoning LLMs in MT are still under-explored.
5.2 Literature Translation
Literature translation is more difficult than news translation since it often requires equivalence beyond the word level Thai et al. (2022). Early literature translation work is limited to small-scale attempts and uses human evaluation to judge model performance Genzel et al. (2010); Jones and Irvine (2013); Besacier and Schwartz (2015); Toral et al. (2018). Recently, Karpinska and Iyyer (2023) utilize GPT-3.5 to perform literature translation, and show that the discourse-level information could enhance model performance. Thai et al. (2022) propose the Par3 dataset, which aims to benchmark LLMs’ literature translation capability from non-English languages to English. Wang et al. (2024) contribute the MetaphorTrans dataset with English-to-Chinese translation. Besides, the sentences in MetaphorTrans generally involve metaphors, thus, it might be suitable for deep reasoning LLMs to perform. DRT Wang et al. (2024) is further proposed based on MetaphorTrans by supervised fine-tuning.
5.3 MT via RL
Machine translation via reinforcement learning has been explored in the literature. Early work tries to train MT models via optimizing BLEU scores Ranzato et al. (2016); Shen et al. (2016); Bahdanau et al. (2017). Wu et al. (2016) design GLEU scores as the rewards to deal with the drawbacks of BLEU in single sentence evaluation. Wu et al. (2018) provide an empirical study on MT via RL, and they propose a method to involve large-scale monolingual data during RL training. Choshen et al. point out the challenges of optimizing MT models via RL, e.g., high-dimensional discrete action space and sparse reward signals. Kiegeland and Kreutzer (2021) provide further analyses on these challenges. Kang et al. (2020) study document-level MT, and they propose a new method to both select context and translate sentences via RL. Though great performance has been achieved, most previous work only focuses on non-reasoning models. After the emergence of deep reasoning LLMs, R1-T1 He et al. (2025) is proposed to optimize deep reasoning LLMs via RL, and it uses COMET scores as the reward signals. However, using COMET as the optimization goal is not well calibrated. In the literature domain, COMET shows low effectiveness as an evaluation metric Karpinska and Iyyer (2023).
6 Conclusion
@@
In this work, we study how to improve the free translation ability of deep reasoning LLMs via RL. We design three types of reward signals in RL: format reward, thought reward and translation reward. Among them, rules and DeepSeek-v3 are used to provide discrete reward values. In this way, using Qwen2.5-7B-Instruct as backbone, we train DeepTrans-7B via GRPO RL algorithm. The experimental results in the literature translation show the effectiveness of DeepTrans-7B. It outperforms QwQ-32B-preview and DeepSeek-Qwen-32B LLMs, and achieves competitive results with DeepSeek-R1 and OpenAI’s o1. In addition, we give deeper analyses on model results during RL training, and find a stylization phenomenon that needs to be carefully considered. Moreover, we summarize the failure experiences and critical components during RL training to provide a deeper understanding of deep reasoning MT LLMs.
Appendix A GRF Evaluator
The GRF prompt borrows from Kocmi and Federmann (2023):
Score the following translation from English to Chinese on a continuous scale from 0 to 100, where score of zero means "no meaning preserved" and score of one hundred means "perfect preservation of meaning, with faithfulness, expressiveness, and elegance". English source: {src} Chinese translation: {hyp} Score:

### v2

@@ 5 Failure Experience
We discuss our failures in reward design, which could also be regarded as the ablation study of $r_{\text{all}}$.
5.1 The Length Penalty of Thought Process.
In the pilot RL experiments, we design a length penalty for the thought content:
$\small r_{\text{length}}=\operatorname{max}(-\frac{\operatorname{len}(\text{think})-\beta}{\eta},0)$ (9)
where $\beta$ and $\eta$ are hyperparameters. $\operatorname{len}\text{(think)}$ indicates the token-level length of the generated thought content. In this manner, the length penalty will punish redundant thinking processes, and we add it to $r_{\text{thought}}$:
$\small r_{\text{thought (w/ penalty)}}=r_{\text{thought}}+r_{\text{length}}$ (10)
We set $\beta$ to 400 and $\eta$ to 400 in the pilot experiments. The DeepTrans-7B model trained with the length penalty is denoted as DeepTrans (w/ LP). As shown in Figure 3, we find that the length penalty does not bring improvement during training. In contrast, it reduces the rate of improvement in model performance. Therefore, we do not adopt the length penalty in the final reward modeling.
@@ 5.2 The Effect of Thought Reward
We also attempt to remove the thought reward, and the variant overall reward is defined as:
$\small r_{\text{all (w/o th)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ r_{\text{trans}}&\text{if }r_{\text{format}}\neq 0\end{cases}$ (11)
However, we find that DeepTrans trained via $r_{\text{all (w/o th)}}$ quickly discards the thought process, and tends to directly output the final translation until the end of the training. To make the model provide a valuable thought process, we further consider a short length penalty on the thought process: If the thought length is less than $k$ tokens, we give a huge penalty to the overall reward:
$\small r_{\text{all}\text{(w/o th \& w/ sp)}}=\begin{cases}0&\text{if }r_{\text{format}}=0\\ 0&\operatorname{len}\text{(think)}<k\\ r_{\text{trans}}&\text{else}\end{cases}$ (12)
We set $k$ to 10, and find that the trained model tends to only provide several empty words in the thought content. For example, “Well, first I have to understand the meaning of the sentence. Then, I will express it in fluent Chinese”. This kind of thought content loses relevance and specificity towards the source sentences.
Besides, the model trained with $r_{\text{all (w/o th)}}$ or $r_{\text{all (w/o th \& w/ sp)}}$ significantly underperforms the original DeepTrans-7B in terms of all metrics. These findings demonstrate the importance of the thought process in model application, and we should consider how to guide models to generate valuable thought during RL training.
@@
w/ 3-point scale w/ 5-point scale w/ 3-point scale w/ 5-point scale
When it was light enough Johnsy, the merciless, commanded that the shade be raised. 当光线足够明亮时,无情的约翰西斯命令把窗帘拉上。 [CometKiwi: 81.72] 当光线已然充足,约翰西,那无情之人,竟令帷幕徐徐卷起[CometKiwi: 74.46] 待天色渐明,那冷漠无情的约翰西,便命人将帘幕徐徐卷起[CometKiwi: 75.08] 天色稍显明亮,约翰西,那颗铁石心肠,便下令将阴影卷起。 [CometKiwi: 66.96] 待天色足够明亮,约翰西这颗铁石心肠,已然下令将帘幕卷起。 [CometKiwi: 71.40]
When one loves one’s Art no service seems too hard. 当一个人热爱自己的艺术时,任何服务都显得不值一提。 [CometKiwi: 70.98] 若钟情于艺道,任劳任怨亦无妨[CometKiwi: 69.38] 一旦钟爱所钟之艺便觉任劳任怨非艰[CometKiwi: 69.38] 艺海倾心无难事,任凭千般苦亦甘[CometKiwi: 54.76] 只要钟爱所钟情的艺术,任何付出都觉轻而易举。 [CometKiwi: 76.59]
For two or three days she went out canvassing for pupils. One evening she came home elated. 连续两三天她出去招生,一个傍晚她兴高采烈地回家了。 [CometKiwi: 83.55] 连续几日外出寻觅门徒一夕归家心欢悦[CometKiwi: 80.25] 她连着几日四处奔走寻觅门徒,直至某晚,她满心欢喜地踏入了家园。 [CometKiwi: 74.54] 连续几日奔走寻门徒,直至傍晚方归家满心欢喜步履轻,笑颜如花映晚霞[CometKiwi: 55.10] 她接连走了两三天,到处奔走寻觅学生。一天傍晚,她满面春风地回到了家中。 [CometKiwi: 83.87]
Table 6: The results of Case Studies on O. Henry.
@@ 5.3 The Effect of Translation Reward
We define the translation reward using a 100-point scale in § 3.1, and we further study the effect of the scoring criteria on translation rewards. Specifically, we use two variants: $r_{\text{trans (3 point)}}$ and $r_{\text{trans (5 point)}}$ simply narrow the scoring scope of $r_{\text{trans}}$ from a 100-point scale to a 3-point scale and a 5-point scale, respectively. During training with these two variants, the $\alpha$ in Eq. 4 is also changed to ensure the same ratio between thought rewards and translation rewards, i.e., we set $\alpha$ to 0.6 and 1.0 when adopting $r_{\text{trans (3 point)}}$ and $r_{\text{trans (5 point)}}$, respectively. We denote the models trained with the variants as “DeepTrans (w/ 3-ps)” and “DeepTrans (w/ 5-ps)”, where “ps” is the abbreviation of “point scale”. As shown in Figure 4 (a), both variants show the sub-optimal results in terms of GRF, verifying the rationality of our original design.
@@
As for CometKiwi in Figure 4 (b), we find that DeepTrans (w/ 5-ps) shows similar trends with original DeepTrans, i.e., first drops and finally recovers. However, DeepTrans (w/ 3-ps) continues to decline and fluctuate during training. In the end, it achieves a low CometKiwi score, and perhaps it needs more training steps to recover the performance. In view of the stylization phenomenon we discussed in § 4.6, we conjecture the trends shown in Figure 4 (b) are attributed to stylization.
@@
To figure it out, we provide case studies on the variant models in Table 6. When the models are trained with 300 steps (1 epoch), both variants show the stylization phenomenon. Further, when trained with 600 steps (2 epochs), we find that DeepTrans (w/ 5-ps) discards the classical-style translations. In contrast, DeepTrans (w/ 3-ps) enhances the classical-style translations. As we also described in § 4.6, the basic reason of the stylization phenomenon is a preference of the reward model, which encourages the MT model to provide stylized translations as a local optimal solution. When the reward model provides a wider scoring scope, it gives a larger latent space for the policy model to explore. Thus, the policy model is more likely to get out of the local optimal solution. If using a narrow reward scope, some minor changes in the policy model will not be reflected by the reward signal immediately, making it hard to transfer the translation strategies. In conclusion, the scoring scope of the reward models has a great effect on the policy model, and it is necessary to employ a wider reward scope for the model to explore a good translation strategy.
6 Conclusion and Future Work
@@
In this work, we propose DeepTrans, which aims to enhance LLMs’ deep reasoning MT ability via RL. We use DeepSeek-v3 as the reward model to provide discrete reward values. The experimental results in the literature MT show the effectiveness of DeepTrans. It outperforms QwQ-32B-preview, and achieves competitive results with state-of-the-art deep reasoning LLMs. In addition, we give deeper analyses on model results during RL training, and find a stylization phenomenon that needs to be carefully considered. Moreover, we summarize the failure experiences and critical components in the RL framework to provide a deeper understanding of deep reasoning MT LLMs.
In the future, the following directions may be worth exploring to promote the deep reasoning translation: (1) Extending deep reasoning translation models into multi-lingual MT. Consequently, the model could think and translate among different languages; (2) Qualifying and controlling the translation style during RL training to alleviate the stylization phenomenon (Section 4.6); (3) Exploring translation rewards with enhanced accuracy, robustness, or efficiency. In this work, we prompt vanilla DeepSeek-v3 to provide the translation reward, which is computationally intensive; (4) Balancing long- and short-thought to improve models’ inference efficiency, and avoid the overthinking issue Sui et al. (2025). For simple translation queries, a long thought process may not be necessary.
7 Limitation
While we show the effectiveness of reinforcement learning in deep reasoning translation, there are some limitations worth noting: (1) We evaluate the deep reasoning translation models in a single translation direction, i.e., English-to-Chinese, and future work could extend our method to other directions; (2) The reward model employed in our experiments is DeepSeek-v3, which requires significant computational resources to infer; (3) Due to resource limitations, we conduct experiments on LLMs with 7B parameters, and future work could extend our method to LLMs with more parameters.